OpenStudy (anonymous):

f(t)= (2t+1)/(t+3) find f '(a) where do you start with this?

OpenStudy (anonymous):

First you need find f'(t) then substitute a on the result.

hartnn (hartnn):

by differentiating f(t) use division u/v rule

hartnn (hartnn):

\(\huge \frac{u'v+uv'}{v^2}\)

OpenStudy (anonymous):

@hartnn \[\frac{ u'v-uv' }{ v^{2} }?\]

hartnn (hartnn):

typo

OpenStudy (anonymous):

could you possibly write what that formula would look like with the functions and derivatives in it?

OpenStudy (anonymous):

For\[f(t) = \frac{ 2t+1 }{ t+3 }\]\[f'(t) = \frac{ \frac{ d(2t+1) }{ dt }\times (t+3) - (2t+1)\times\frac{ d(t+3) }{ dt }}{ (t+3)^{2} }\]

OpenStudy (anonymous):

@darkmare I hope it helped.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are welcomed!